# WaveStitch: Flexible and Fast Conditional Time Series Generation with Diffusion Models

### v2

@@ 2.4. Research Gaps
The inability to generalise to unseen conditional patterns limits training time methods. For instance, TimeWeaver (Narasimhan et al., 2024) conditions only on metadata and cannot adapt based on observed signals. Similarly, SSSD (Alcaraz and Strodthoff, 2023) requires knowing the conditional mask of observations, which can lead to overfitting on the patterns seen during training. Classifier-free guidance requires both the conditional and unconditional model variants, which still requires training the conditional branch with specific conditions in mind (Ho and Salimans, 2021). Even classifier guidance depends on pre-trained classifiers tailored to each condition type (Liu et al., 2024). Hence, it is not flexible during inference since the classifiers require training with the right conditions in advance.
@@ 3.1. Overview
We first train an metadata-conditioned denoiser $f_{\theta}$ to predict noise using only metadata, without relying on specific patterns of observed signals. This approach allows WaveStitch to generate reasonable signal estimates even in the worst-case scenario with just the metadata and no observed signals. WaveStitch then iteratively refines the generated samples during inference via a gradient-based correction. At each denoising step, we adjust the generated sample using the gradient of a novel conditional loss computed over observed signals. This approach is similar to gradient inversion attacks (Jeon et al., 2021), where an attacker iteratively optimises a dummy input using gradients to retrieve the original data while keeping the model frozen.
@@ 3.2. Preprocessing Categorical Metadata
Time series metadata often contains recurring categorical attributes, such as day-of-the-week, month, hour, etc. To capture the periodic structure in these features, we use sine and cosine transformations inspired by positional encodings in Transformers (Vaswani, 2017) and related work in time series generation (Suh et al., 2024). For a given feature, we map each category $k\in\{0,1,\dots,K-1\}$ to an angle $\theta_{k}=\frac{2\pi k}{K}$ on the unit circle. Each categorical value $k$ is then encoded by the coordinates $(\sin(\theta_{k}),\cos(\theta_{k}))$, reflecting the feature’s cyclical nature. This representation is more compact than one-hot encoding  (Rodr´ıguez et al., 2018; Kotelnikov et al., 2023) and represents periodic behaviour more naturally. This step is done as a precursor to both the training and inference.
@@
3.3. Training Metadata-conditioned Denoiser
We illustrate our training process in Figure 5 and show the steps in Algorithm 1. It begins with dividing the entire dataset sequence of $N$ timesteps into overlapping windows of size $w$ (line 4), using a stride of one. Then, we noise the window up to a randomly sampled diffusion step $t$ and estimate the added noise, following the standard DDPM training process (Ho et al., 2020) (lines 6,7). The noise is estimated by conditioning on the metadata for the window, $\mathbf{a}^{(i:i+w-1)}$, and uses the MSE loss to update $f_{\theta}$’s parameters (line 8). This way, the model learns the interactions between the metadata and the signals without relying on any observed signals. This separation sets the foundation to flexibly condition on the observations during inference (challenge 2). As we showed in 4(c), leveraging the observed values can further improve generation quality, particularly for fine-grained tasks where observed signals from neighbouring timesteps can yield higher precision.
Figure 5. Metadata-conditioned Denoiser Training
@@ 3.4. Signal-conditioned Inference
Our process for conditional generation at inference consists of several steps, which we detail as follows:
@@
We define the test set $\mathcal{X}=\{\mathbf{x}^{(i)}\}_{i=1}^{M}$, where $\mathbf{x}^{(i)}$ indicates the observed signals, and the unknown values are left empty, representing the portions to be generated. The complete set of metadata, $\mathcal{A}=\{\mathbf{a}^{(i)}\}_{i=1}^{M}$ is available as part of the setup, consistent with the problem definition (Sec. 2.3). A corresponding mask $\mathcal{M}=\{\mathbf{m}^{(i)}\}_{i=1}^{M}$ is defined, where $\mathbf{m}^{(i)}=1$ indicates that $\mathbf{x}^{(i)}$ requires generation and $\mathbf{m}^{(i)}=0$ marks the observations. We first divide the test set into overlapping time windows of size $w$, spaced by a stride $s$, yielding $\mathcal{X}_{w}=\{\mathbf{x}_{w}^{(j)}\}_{j=1}^{\left\lfloor\frac{M-w}{s}\right\rfloor+1}$, $\mathcal{M}_{w}=\{\mathbf{m}_{w}^{(j)}\}_{j=1}^{\left\lfloor\frac{M-w}{s}\right\rfloor+1}$, and $\mathcal{A}_{w}=\{\mathbf{a}_{w}^{(j)}\}_{j=1}^{\left\lfloor\frac{M-w}{s}\right\rfloor+1}$. Following this, we proceed with the generation process detailed in Algorithm 2, which iterates through steps of $t=T,T-1\dots 1$. At each step, the metadata-conditioned denoiser first produces an intermediate sample, which is then adjusted using the conditional loss gradients.
@@
We first initialise the signal to be generated ($\mathcal{\hat{X}}_{w}$) in line 2. Then we divide $\mathcal{X}_{w}$, $\mathcal{M}_{w}$, and $\mathcal{A}_{w}$ into mini-batches of size $b$, each containing samples of $w$ timesteps (line 3). To condition on available signals and metadata, we first get a dirty estimate of the fully-denoised signal ($\mathbf{\hat{x}}_{w,0}^{(j)}$) using our metadata-conditioned denoiser. We get this estimate by reversing the fast-forward noising step Eq. (2) as follows:
@@
Here, $\mathbf{\hat{x}}_{w,t}^{(j)}$ is the $j\textsuperscript{th}$ window’s estimate at the $t\textsuperscript{th}$ diffusion step.
@@
The denoiser $f_{\theta}$ estimates samples by only conditioning on the metadata. To further factor in the observed signal values, we correct the samples using the gradients from the conditional loss, defined as follows:
@@
Self-Guidance: The self-guidance loss (Kollovieh et al., 2024) ensures the generated signal values align with the observed signals within a given time window. It is defined as follows:
@@
Stitching: Since self-guidance treats each window independently, it may produce inconsistencies across window boundaries by ignoring overlaps. These overlaps are critical for maintaining temporal continuity, and failing to align them can lead to incoherent transitions. Stitching overcomes this limitation by enforcing consistency in overlapping regions, restricting the solution space to a more realistic subset than those produced by self-guidance alone. For a given window $\mathbf{\hat{x}}_{w,t}^{(j)}$, we first get a denoised estimate $\mathbf{\hat{x^{\prime}}}_{w,t-1}^{(j)}$ using the standard reverse diffusion step Eq. (3) without any conditioning:
@@
We then ”stitch” these uncorrected estimates across window boundaries to ensure coherence. Specifically, for a window of size $w$ and stride $s$, the overlap consists of the first wsw-s timesteps of $\mathbf{\hat{x^{\prime}}}_{w,t-1}^{(j)}$ and the last wsw-s timesteps of $\mathbf{\hat{x^{\prime}}}_{w,t-1}^{(j-1)}$. The stitch loss then penalises discrepancies in this overlap as follows [^3] For the first window, we define $\mathcal{L}_{\text{stitch}}^{(1)}=0$, as there are no preceding windows.:
@@
Adjustment: To dynamically condition on the observed signals, we first denoise one step Eq. (7). Then, we adjust it using the gradient scaled by guidance coefficient $\eta$ as follows:
(9) $\mathbf{\hat{x}}_{w,t-1}^{(j)}=\mathbf{\hat{x^{\prime}}}_{w,t-1}^{(j)}-\eta\nabla_{\mathbf{\hat{x^{\prime}}}_{w,t-1}^{(j)}}\mathcal{L}_{\text{cond}}^{(j)},$
where $\mathcal{L}_{\text{cond}}$ is as defined in Eq. (5), $z\sim\mathcal{N}(0,I)$, and $\sigma_{t}=(1-\alpha_{t})\times\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}$. At the end of the denoising process, we reintroduce the observations (line 11) and merge the sliding windows (line 12), to yield the final sequence. Figure 6 illustrates the conditional denoising process using the combination of the two losses.
To summarize, Eq. (7) computes an initial denoised estimate without any conditioning. If left unmodified, the reverse diffusion trajectory would follow this path, ignoring both observations and inter-window coherence. Equation (9) adjusts the estimate by descending the gradient of the conditional loss, which consists of two terms. The self-guidance loss Eq. (6) enforces alignment with the observed signal values within each window, while the stitch loss Eq. (8) enforces consistency across overlapping windows. Together, these components nudge the model at every denoising step to yield samples that adhere to observations and remain globally coherent. The following theoretical analysis (Sec. 3.5) formalises how minimising Eq. (5) restricts the generated samples to a subset that is consistent with observations and coherent across windows, to justify why this conditioning strategy effectively guides the diffusion process toward realistic and temporally consistent time series.
@@ 3.5. Theoretical Analysis
Here, we formalise how the inference-time conditional loss, comprising self-guidance and stitching, constrains generation towards a tighter solution space that promotes coherence and observation-alignment while remaining within the training distribution.
Definitions. Let $\mathcal{X}_{\text{cond}}^{\delta}$} be the set of samples for which the conditional loss is at most $\delta$ $\geq 0$, i.e., they respect both the overlap consistencies and the observed signals within windows up to a tolerance $\delta$. As a special case, $\mathcal{X}_{\text{cond}}^{0}$} represents the set of samples with zero-conditional loss. Similarly, let $\mathcal{X}_{\text{stitch}}^{\delta}$} and $\mathcal{X}_{\text{self}}^{\delta}$} be the set of samples with at most $\delta$ stitch and self-guidance losses respectively. Finally, we define $\mathcal{X}_{\text{real}}$ as the set of real samples, which must obey the following conditions: (a) exactly match the known values, i.e., zero self-guidance loss; and (b) be temporally coherent, i.e., any overlapping windows yield identical values on their shared timesteps (zero stitch loss). A violation of these conditions would result in conflicting predictions at the same timestep upon merging, which is not possible for a real time series.
@@ 4. Pipelined Parallel Design of WaveStitch
Algorithm 2 Parallel Denoising with Inference Conditioning
2:Initialize outputs:
$\mathcal{\hat{X}}_{w}=\{\mathbf{\hat{x}}_{w,T}^{(j)}\sim\mathcal{N}(0,I)\}_{j=1}^{\left\lfloor{(M-w)}/{s}\right\rfloor+1}$
4:for each mini-batch do
5:  for ($\mathbf{x}_{w}^{(j)},\mathbf{a}_{w}^{(j)},\mathbf{m}_{w}^{(j)}$) in mini-batch in paralleldo
6:   for step $t=T,T-1,\dots,1$ do
7:     Initial estimate $\mathbf{\hat{x}}_{w,0}^{(j)}$ using Eq. (4)
9:     Compute conditional loss $\mathcal{L}_{\text{cond}}^{(j)}$ using Eq. (5)
11:   Re-introduce observations:
12:Merge windows:
13:return $\mathcal{\hat{X}}$
The conditioning mechanism described in Sec. 3.4 could be applied autoregressively by processing each window sequentially. Most steps in Algorithm 2 can be performed independently across windows. We additionally make the conditional loss computation (Algorithm 2, line 9) parallel-compatible by a parallel shift (or roll) operation to align overlapping regions across all windows, to compute the loss between pairs of windows concurrently. This process resembles a pipeline, as shown in Figure 6. Although dependencies are enforced only with the immediately preceding window, coherence emerges gradually over successive denoising iterations. As dependencies propagate through the pipeline, non-overlapping parts adjust to maintain coherence with their neighbours.
The stride $s$ is crucial in balancing efficiency and coherence. When $s$ is large enough to eliminate overlap, the process becomes a divide-and-conquer strategy (Li et al., 2024; Yin et al., 2023), that synthesises windows independently. Conversely, smaller strides enforce stronger dependencies between windows due to larger overlapping regions. In the extreme case of $s=1$, maximum coherence is achieved as nearly all timesteps are shared between consecutive windows. However, the optimal stride depends on the time series itself. Smaller strides ensure better alignment for signals with strong temporal dependencies, while larger strides may suffice for weaker ones. This flexibility allows our method to adapt to varying degrees of temporal dependence for efficient and coherent signal generation.
Time Complexity and Speedup: WaveStitch substantially improves the generation speed compared to autoregression due to parallelism. Using the notation in Algorithm 2, each dataset has around $\frac{M-w}{s}$ time windows, where $M$ is the total length, $w$ is the window size, and $s$ is the stride. Since autoregression processes each window sequentially, the total cost is $\mathcal{O}(T\cdot t_{\theta}(w)\cdot(M-w)/s)$, where $t_{\theta}(w)$ is the cost for denoising a window by one step. In contrast, the parallel approach processes $b$ windows in parallel per mini-batch, which significantly reduces the time complexity to $\mathcal{O}(T\cdot t_{\theta}(w)\cdot\frac{M-w}{b\cdot s})$. This analysis suggests an ideal speedup factor of approximately $b$ over the autoregressive approach, representing an upper bound on performance optimal hardware utilisation.

### v3

@@ 2.4. Research Gaps
The inability to generalise to unseen conditional patterns limits training time methods. For instance, TimeWeaver (Narasimhan et al., 2024) conditions only on the metadata and cannot leverage information from the observed signals. Similarly, SSSD (Alcaraz and Strodthoff, 2023) requires knowing the conditional mask of observations, which can lead to overfitting on the patterns seen during training. Classifier-free guidance requires both a conditional and unconditional model variant, which still requires training the conditional one with specific conditions in mind (Ho and Salimans, 2021). Even classifier guidance depends on pre-trained classifiers tailored to each condition type (Liu et al., 2024), making it inflexible during inference since the classifiers require training with the right conditions in advance.
@@ 3.1. Overview
We first train a metadata-conditioned denoiser $f_{\theta}$ to predict noise using only metadata, without relying on specific patterns of observed signals. This approach enables WaveStitch to generate reasonable signal estimates even in the worst-case scenario, where we are given only the metadata and no observed signals as conditions. WaveStitch then iteratively refines the generated samples during inference via a gradient-based correction. At each denoising step, we adjust the generated sample using the gradient of a novel conditional loss computed over observed signals. This approach is similar to gradient inversion attacks (Jeon et al., 2021), where an attacker iteratively optimises a dummy input using gradients to retrieve the original data while keeping the model frozen.
@@ 3.2. Preprocessing Categorical Metadata
Time series metadata often contains recurring categorical attributes, such as day of the week, month, and hour. To capture the periodic structure in these features, we use sine and cosine transformations inspired by positional encodings in Transformers (Vaswani, 2017) and related work in time series generation (Suh et al., 2024). For a given feature, we map each category $k\in\{0,1,\dots,K-1\}$ to an angle $\theta_{k}=\frac{2\pi k}{K}$ on the unit circle. Each categorical value $k$ is then encoded by the coordinates $(\sin(\theta_{k}),\cos(\theta_{k}))$, reflecting the feature’s cyclical nature. This representation is more compact than one-hot encoding  (Rodríguez et al., 2018; Kotelnikov et al., 2023) and represents periodic behaviour more naturally.
3.3. Training Metadata-conditioned Denoiser
We illustrate our training process in Figure 5 and show the steps in Algorithm 1. It begins with dividing the entire dataset with $N$ timesteps into overlapping windows of size $w$ (line 4), using a stride of one. We then forward the noise up to a randomly sampled diffusion step $t$ and estimate the added noise, following the standard DDPM training process (Ho et al., 2020) (lines 6, 7). The noise is estimated by conditioning on the metadata for the window, $\mathbf{a}^{(i:i+w-1)}$, and uses the MSE loss to update $f_{\theta}$’s parameters (line 8). This way, the model learns the interactions between the metadata and the signals without relying on any observed signals. This separation sets the foundation to flexibly condition on the observations during inference (challenge 2). As we showed in 4(c), leveraging the observed values can further improve generation quality, particularly for fine-grained tasks where the observed signals from neighbouring timesteps can yield higher precision.
@@ 3.3. Training Metadata-conditioned Denoiser
Figure 5. Metadata-conditioned denoiser training.
@@ 3.4. Signal-conditioned Inference
With our metadata-conditioned denoiser (Sec. 3.3), we additionally condition on the observed signals that are injected directly at inference, which we detail as follows:
@@
We define the test set $\mathcal{X}=\{\mathbf{x}^{(i)}\}_{i=1}^{M}$, where $\mathbf{x}^{(i)}$ indicates the observed signals, and the unknown values are left empty, representing the portions to be generated. The complete set of metadata, $\mathcal{A}=\{\mathbf{a}^{(i)}\}_{i=1}^{M}$ is available as part of the setup, consistent with the problem definition (Sec. 2.3). A corresponding mask $\mathcal{M}=\{\mathbf{m}^{(i)}\}_{i=1}^{M}$ is defined, where $\mathbf{m}^{(i)}=1$ indicates that $\mathbf{x}^{(i)}$ requires generation and $\mathbf{m}^{(i)}=0$ indicates that the corresponding signals are observed. We first divide the test set into overlapping time windows of size $w$, spaced by a stride $s$, yielding $\mathcal{X}_{w}=\{\mathbf{x}_{w}^{(j)}\}_{j=1}^{\left\lceil\frac{M-w}{s}\right\rceil+1}$, $\mathcal{M}_{w}=\{\mathbf{m}_{w}^{(j)}\}_{j=1}^{\left\lceil\frac{M-w}{s}\right\rceil+1}$, and $\mathcal{A}_{w}=\{\mathbf{a}_{w}^{(j)}\}_{j=1}^{\left\lceil\frac{M-w}{s}\right\rceil+1}$. Following this, we proceed with the generation process detailed in Algorithm 2, which iterates through steps of $t=T,T-1\dots 1$. At each step, the metadata-conditioned denoiser first produces an intermediate sample, which is then adjusted using conditional loss gradients.
Algorithm 2 Parallel Denoising with Inference Conditioning
2:Initialize outputs:
$\mathcal{\hat{X}}_{w}=\{\mathbf{\hat{x}}_{w,T}^{(j)}\sim\mathcal{N}(0,I)\}_{j=1}^{\left\lceil{(M-w)}/{s}\right\rceil+1}$
4:for each mini-batch do
5:  for ($\mathbf{x}_{w}^{(j)},\mathbf{a}_{w}^{(j)},\mathbf{m}_{w}^{(j)}$) in mini-batch in parallel: do
6:   for step $t=T,T-1,\dots,1$ do
7:     Dirty estimate $\mathbf{\hat{x}}_{w,0}^{(j)}$ using eq. (4)
9:     Compute conditional loss $\mathcal{L}_{\text{cond}}^{(j)}$ using eq. (5)
11:   Re-introduce observations:
12:Merge windows:
13:return $\mathcal{\hat{X}}$
@@
We first initialise the signal to be generated ($\mathcal{\hat{X}}_{w}$) in line 2. Then we divide $\mathcal{X}_{w}$, $\mathcal{M}_{w}$, and $\mathcal{A}_{w}$ into mini-batches of size $b$, each containing samples of $w$ timesteps (line 3). To condition on the observed signals and metadata, we first get a dirty estimate of the fully-denoised signal ($\mathbf{\hat{x}}_{w,0}^{(j)}$) using our metadata-conditioned denoiser (line 7). We get this estimate by reversing the fast-forward noising step in eq. (2) as follows:
@@
Here, $\mathbf{\hat{x}}_{w,t}^{(j)}$ is the model’s estimate at the $t\textsuperscript{th}$ diffusion step.
@@
The denoiser $f_{\theta}$ estimates samples by only conditioning on the metadata. To further factor in the observed signal values, we correct the samples using the gradients from the conditional loss (line 9), defined as follows:
@@
Self-Guidance: The self-guidance loss (Kollovieh et al., 2024) ensures that the generated outputs match the observed signals within a given time window. It is defined as follows:
@@
Stitching: Since self-guidance treats each window independently, it may produce inconsistencies across window boundaries by ignoring overlaps. These overlaps are crucial for maintaining temporal continuity, and failing to align them can result in incoherent transitions. Stitching overcomes this limitation by enforcing consistency in overlapping regions, restricting the solution space to a more realistic subset than those produced by self-guidance alone. For a given window $\mathbf{\hat{x}}_{w,t}^{(j)}$, we first denoise one step $\mathbf{\hat{x^{\prime}}}_{w,t-1}^{(j)}$ unconditionally using eq. (3) (line 8) :
@@
We then “stitch” these uncorrected estimates across window boundaries to ensure coherence. Specifically, for a window of size $w$ and stride $s$, the overlap consists of the first wsw-s timesteps of $\mathbf{\hat{x^{\prime}}}_{w,t-1}^{(j)}$ and the last wsw-s timesteps of $\mathbf{\hat{x^{\prime}}}_{w,t-1}^{(j-1)}$. The stitch loss then penalises discrepancies in this overlap as follows [^3] For the first window, we define $\mathcal{L}_{\text{stitch}}^{(1)}=0$, as there are no preceding windows.:
@@
Adjustment: To dynamically condition on the observed signals, we correct the unconditional outputs from eq. (7) using the conditional loss gradients scaled by guidance coefficient $\eta$ (line 10):
(9) $\mathbf{\hat{x}}_{w,t-1}^{(j)}=\mathbf{\hat{x^{\prime}}}_{w,t-1}^{(j)}-\eta\nabla_{\mathbf{\hat{x}}_{w,t}^{(j)}}\mathcal{L}_{\text{cond}}^{(j)},$
where $\mathcal{L}_{\text{cond}}$ is as defined in eq. (5), $z\sim\mathcal{N}(0,I)$, and $\sigma_{t}=(1-\alpha_{t})\times\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}$. At the end of the denoising process, we reintroduce the observations (line 11) and merge the overlapping windows (line 12), to yield the final sequence. Figure 6 illustrates the conditional denoising process using the combination of self-guidance and stitching.
To summarise, eq. (7) unconditionally denoises the sample without enforcing any constraints. If left unmodified, the reverse diffusion trajectory would follow this path, ignoring both observations and inter-window coherence. Equation (9) adjusts the trajectory using the gradient of the conditional loss, which consists of two terms. The self-guidance loss eq. (6) enforces alignment with the observed signal values within each window, while the stitching eq. (8) enforces consistency across overlapping windows. Together, these components nudge the model at every denoising step to yield samples that adhere to observations and remain globally coherent.
@@ 3.5. Theoretical Analysis
Here, we formalise how the inference-time conditional loss, comprising self-guidance and stitching, constrains generation towards a tighter solution space that promotes global temporal coherence and alignment with the given signal observations.
Definitions. Let $\mathcal{X}_{\text{cond}}^{\delta}$} be the set of samples for which the conditional loss is at most $\delta$ $\geq 0$, i.e., they respect both the overlap consistencies and the observed signals within windows up to a tolerance $\delta$. As a special case, $\mathcal{X}_{\text{cond}}^{0}$} represents the set of samples with zero-conditional loss. Similarly, let $\mathcal{X}_{\text{stitch}}^{\delta}$} and $\mathcal{X}_{\text{self}}^{\delta}$} be the set of samples with at most $\delta$ stitch and self-guidance losses respectively. Finally, we define $\mathcal{X}_{\text{real}}$ as the set of real samples, which must obey the following conditions: (a) exactly match the known values, i.e., zero self-guidance loss; and (b) be temporally coherent, i.e., any overlapping windows yield identical values on their shared timesteps (zero stitch loss). A violation of condition (b) would result in conflicting predictions at the same timestep upon merging, which is not possible for a real time series.
@@ 4. Pipelined Parallel Design of WaveStitch
The conditioning mechanism described in Sec. 3.4 could be applied autoregressively by processing each window sequentially. However, most steps in Algorithm 2 can be performed independently across windows. Even the conditional loss computation (Algorithm 2, line 9) can be made parallel-compatible by a parallel shift (or roll) operation to align overlapping regions, to compute the pairwise loss between consecutive windows concurrently. This process resembles a pipeline, as shown in Figure 6. At each denoising step, non-overlapping parts adjust to maintain coherence with their neighbours. Although these adjustments are only enforced with the neighbouring window, coherence gradually propagates over the whole pipeline over successive steps.
The stride $s$ is crucial in balancing efficiency and coherence. When $s$ is large enough to eliminate overlap, the process becomes a divide-and-conquer strategy (Li et al., 2024; Yin et al., 2023), that synthesises windows independently. Conversely, smaller strides enforce stronger dependencies between windows due to larger overlaps. In the extreme case of $s=1$, maximum coherence is achieved as nearly all timesteps are shared between consecutive windows. However, the optimal stride depends on the time series itself. Smaller strides ensure better alignment for signals with strong temporal dependencies, while larger strides may suffice for weaker ones. This flexibility allows our method to adapt to varying degrees of temporal dependence for efficient and coherent signal generation.
Time Complexity and Speedup: WaveStitch substantially improves the generation speed compared to autoregression due to parallelism. Using the notation in Algorithm 2, each dataset has around $\frac{M-w}{s}$ time windows, where $M$ is the total length, $w$ is the window size, and $s$ is the stride. With sequential execution, the total cost is $\mathcal{O}(T\cdot t_{\theta}(w)\cdot(M-w)/s)$, where $t_{\theta}(w)$ is the time taken to denoise a window by one step. In contrast, the parallel approach processes $b$ windows in parallel per mini-batch, which significantly reduces the time complexity to $\mathcal{O}(T\cdot t_{\theta}(w)\cdot\frac{M-w}{b\cdot s})$. This analysis suggests an ideal speedup factor of approximately $b$ over the autoregressive approach, representing an upper bound on performance under optimal hardware utilisation.